OpenStudy (anonymous):

a 20 kg projectile has a velocity og 600 m/s. the shell acquired the velocity in a mortar (barrel) 3m long. Determine the magnitude of the average force against the shell as it was fired.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

First Find the acceleration

OpenStudy (anonymous):

V^2 - u^2 = 2as

OpenStudy (anonymous):

a=?

OpenStudy (anonymous):

find....

OpenStudy (anonymous):

a= V-Vo/t

OpenStudy (anonymous):

we r not given time lol

OpenStudy (anonymous):

so use this V^2 - u^2 = 2as

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

1.2 * 10^5

OpenStudy (anonymous):

wat is ur answer @experimentX

OpenStudy (anonymous):

I FOUNDED

OpenStudy (anonymous):

icant get it

OpenStudy (anonymous):

Do u have answer key

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wew

OpenStudy (anonymous):

its an example in my book and i dont know how to anwer

OpenStudy (shane_b):

Maybe seeing it all worked in one post will help: \[V_i=0m/s\]\[V_f=600m/s\]\[s=3m\]\[m=20kg\]\[V_f^2-V_i^2=2as\]\[(600m/s)^2-(0m/s)^2)=2(a)(3m)\]\[a=60,000 m/s^2\]\[F=ma=(20kg)(60,000m/s^2)=1.2x10^6N\]